Sara’s Salsa Company produces its condiments in two types: Extra Fine for restaurant customers and Family Style for home use. Salsa is prepared in department 1 and packaged in department 2. The activities, overhead costs, and drivers associated with these two manufacturing processes and the company’s production support activities follow.

Process Activity Overhead cost Driver Quantity
Department 1 Mixing $ 5,300 Machine hours 1,700
Cooking 10,000 Machine hours 1,700
Product testing 113,300 Batches 550
$ 128,600
Department 2 Machine calibration $ 290,000 Production runs 800
Labeling 16,000 Cases of output 130,000
Defects 10,000 Cases of output 130,000
$ 316,000
Support Recipe formulation $ 98,000 Focus groups 50
Heat, lights, and water 34,000 Machine hours 1,700
Materials handling 73,000 Container types 10
$ 205,000


Additional production information about its two product lines follows.

Extra Fine Family Style
Units produced 28,000 cases 102,000 cases
Batches 280 batches 270 batches
Machine hours 650 MH 1,050 MH
Focus groups 20 groups 30 groups
Container types 7 containers 3 containers
Production runs 280 runs 520 runs

rev: 06_09_2018_QC_CS-128873

Required:
1. Using a plantwide overhead rate based on cases, compute the overhead cost that is assigned to each case of Extra Fine Salsa and each case of Family Style Salsa.
2. Using the plantwide overhead rate, determine the total cost per case for the two products if the direct materials and direct labor cost is $4 per case of Extra Fine and $3 per case of Family Style.
3.a. If the market price of Extra Fine Salsa is $13 per case and the market price of Family Style Salsa is $7 per case, determine the gross profit per case for each product.
3.b. What might management conclude about the Family Style Salsa product line?
Answer #1

ANSWER

1)

Total overhead cost = 128,600+316,000+205,000=649,600

Total cases = 28,000+102,000=130,000

overhead cost per caset =Total overhead /total case

                                     = 649,600/ 130,000

                                      = $ 4.99 per case

2)

Extra cost = 4 +4.99 = 8.99 per case

Family = 3+4.99 = 7.99 per case

3.a)

Gross profit for extra =Price -cost

                                  = 13 - 8.99 = 4.01 per case

For family = 7 -7.99 = - .99 per case

3.b)

The management should eliminate family style as it results in negative gross profit of $ -.99 per case

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************